OpenStudy (goformit100):

What is the differentiation of :- |-7| + 3|4|

OpenStudy (yttrium):

It's automatically zero.

hartnn (hartnn):

because |-7| + 3|4| is constant

OpenStudy (yttrium):

But let say it is |-7x|, how can we derive that @hartnn

OpenStudy (yttrium):

I think it's just 7x.

hartnn (hartnn):

derivative of |x| is |x|/x or x/|x|

OpenStudy (yttrium):

Oww. New learning. Thanks @hartnn . So it can be 7|x|/x, right?

hartnn (hartnn):

yep correct. the other equivalent way is to use \(|a| = \sqrt {a^2} \) and that would give you 7x/sqrt (x^2) which is same thing :)

OpenStudy (anonymous):

12 You guys are going too complex its just simple math

hartnn (hartnn):

how is it 12 ?

OpenStudy (anonymous):

|-7| is absolute value which is how far from zero which is 7

OpenStudy (anonymous):

the rest is just 3 times 4

hartnn (hartnn):

yep |-7| + 3|4| = 7+12 =19 but he asked for the differentiation :)

OpenStudy (goformit100):

Are there two cases for it.

OpenStudy (anonymous):

i keep getting it wrong. LOL

hartnn (hartnn):

2 but equivalent cases

OpenStudy (anonymous):

my bad your wright its 19 i just thought it was 12 beacuse 3 x 4 =12